Calculate the pH of an aqueous solution of 0.571 M trimethylamine (a weak base with the formula (CH3)3N) . The value of Kb is 6.30×10-5.

pH =
1
Expert's answer
2020-10-30T04:02:14-0400

trimethylamine has the formula

= (CH3)N

It is a mono acidic base.

Concentration of trimethylamine

= 0.571 M

Base constant of , Kb = 6.30×10–5

We know,

pOH = 0.5(pKb – logC)

= 0.5(pKb – logC)

= 0.5(–logKb – logC)

= 0.5[–log(6.30×10–5)– log0.571]

= 2.222


Again, pKw = pOH + pH

Or, pH = pKw – pOH

Or, pH = 14 – 2.222

(at 25°C pKw = 14)

Or, pH = 11.778



Hence,

pH of given trimethylamine is

= 11.778